What is the most likely identity of the species undergoing titration?

13 12 10 8 pH 7- 4 2 10 30 50 70 90 110

a) hydrobromic acid (HBr); Ka = 1.08 x 109

b) acetic acid (CH3CO2H); Ka = 1.8 x 10-5

c) ammonia (NH3); Kb = 1.8 x 10-5

d) sodium acetate (NaC2H3O2); Kb = 5.6 x 10-10
Answer #1

pH is equal to pKa at the center of vertical part of curve

SO, pKa here is 9

use:
pKa = -log Ka
9 = -log Ka
Ka = 1.0*10^-9

Kb = 1*10^-14/Ka
= 1*10^-14/(1*10^-9)
= 1*10^-5

This is closest to kb of NH3

Answer: c